Determine the breakeven resale price 10 years from now of an apartment house that can be bought today for $449,000. Its annual net income is $54,000. The owner wants a10% annual return on her investment.
Answer #1

Return on investment = 10 % = 0.1

initial investment = 449000

annual net income = 54000

no.of years = 10

Break even resale price = inital investment * ( 1 + annual return of investment)^10 - (net annual income) * ( 1 + 1.1 + 1.1^2 + 1.1^3 .... +1.1^9)

Resale price = 449000* (1+0.1)^10 + 54000 + (1.1^10 -1)/0.1 =$1,218,606.3

answered by: Benn
Answer #2

F=449000(F/P, 10%, 10) + 54000(F/A, 10%, 10)

= 2,025,304

 

DZ
Answer #3

F=54000(F/A, 10%, 10) - 449000(F/P, 10%, 10)

= -304216

 

Missing that amount to break even at a 10% interest rate

 

DZ
Answer #4

Is their an answer to this question ? I am getting 304108.
Answer #5

449000(F/P,10%,10)  = 54000 (F/A,10%,10) + Resale Price

Solve for Resale Price and get 304108
Answer #6

test